OpenStudy (anonymous):

Solve the Liner system x-4y=10 2x+y=2

OpenStudy (anonymous):

@uri Hate to ask you again but this is almost the same as the question you just helped me with :) these liner equations are driving me crazy!

OpenStudy (uri):

It's easy,Okay a bit different.I'll help ya We basically have to make the value of y same but different signs so it can be cancelled So, x-4y=10 2x+y=2(multiply by 4) x-4y=10 8x+4y=8 ====== 9x=18 x=18/9 x=2 Now find y using the ORIGINAL eq 2 and putting the value of x @JoStudy

OpenStudy (anonymous):

OK where did you get the multiply by 4? and let's see if I can solve for Y 2(2)+y=2 4+y=2 4-4+y=2-4 y=-2 So the liner solution would be (2,-2)

OpenStudy (anonymous):

Is your profile picture totally white or is it my computer?

OpenStudy (uri):

It's common sense xD If we want to make x as 4x we muliply by 4 duh and yea you did y right :) Correct ha!

OpenStudy (uri):

My pic is totally white..:3 i always don't like putting up my face c:

OpenStudy (anonymous):

OK! So we multiply everything by 4! So to make x 4x we have to multiply which means we have to multiply everything else! *lightbulb goes off*

OpenStudy (uri):

Yes smarty :3

OpenStudy (anonymous):

Thanks @uri :)